# On the extraction of Alcock-Paczyński signal from voids: a novel approach via reconstruction

### v1

@@
On the extraction of Alcock-Paczyński signal from voids:
@@
The void-galaxy cross-correlation function is a powerful tool to extract cosmological information. Through the void-galaxy cross-correlation function, cosmic voids, the underdense regions in the galaxy distribution, are used for refined deductions of the Universe’s content by correcting apparent geometric distortions. This study proposes a novel procedure for optimally extracting the Alcock-Paczyǹski (AP) signal from cosmic voids through a cosmological reconstruction technique. Employing cosmological reconstruction, specifically using the Zel’dovich approximation, we estimate the true positions of galaxies from their redshift-space locations, reducing distortions introduced by peculiar velocities. Unlike previous analyses, we identify voids and measure the void–galaxy cross-correlation function directly in reconstructed space. This approach enables us, for the first time, to include in our analysis small nonlinear voids, typically discarded in previous studies, thus enhancing the statistical power of void studies and significantly improving their cosmological constraining power. Reconstruction is particularly effective even at small scales for voids, due to their clean and dynamically simple environment. This ability to recover information encoded on small scales significantly enhances the precision of the analysis, leading to a $\sim 23\%$ improvement in the constraints on the AP parameters compared to previous methods where the analysis is performed in redshift space and, consequently, to a better estimate of the derived cosmological parameters. Our analysis also includes a comprehensive set of consistency checks, demonstrating its robustness. We expect this methodology to yield a substantial gain in constraining power when applied to data from modern large-scale structure surveys.
@@ 1 Introduction
Cosmic voids, vast underdense regions that populate the large-scale structure of the Universe, offer a unique and insightful laboratory for cosmology (Pisani et al. 2019; Kreisch et al. 2019; Moresco et al. 2022). Due to the statistically isotropic nature of structure formation, these initially irregular underdensities evolve toward an approximately spherical shape as matter evacuates their interiors under gravitational dynamics. Although individual voids deviate from sphericity, the process of rescaling by the effective radii and stacking multiple voids yields an average spherically symmetric profile that robustly characterizes their global properties. This averaged spherical structure provides a powerful tool for cosmological analysis. In particular, it enables the application of the Alcock-Paczyǹski (AP) test (Alcock & Paczynski 1979), which relies on the fact that the observed void profile is obtained by converting galaxy positions observed through directly measurable quantities, i.e., angular coordinates ($\mathbf{\Theta}=(\theta,\phi)$), and redshift ($z$), into comoving distances under an assumed cosmological model. If the fiducial model differs from the true one, the inferred galaxy distribution is systematically distorted, breaking the isotropy of stacked voids. By measuring such distortions in the stacked void profile, the AP test allows us to extract valuable information about the expansion history and geometry of the Universe, as well as the dark energy equation of state.
@@
The idea of applying the AP test to cosmic voids was first proposed by Ryden (1995), although its implementation on individual voids proved challenging due to their complex and irregular shapes. This limitation was addressed by Lavaux & Wandelt (2012), who introduced the concept of stacking voids to average out shape irregularities and enhance the statistical power of the test. The first practical applications of the AP test to stacked voids were carried out by Sutter et al. (2012) and Mao et al. (2017). These developments have led to a shift in methodology, favoring the analysis of AP distortions with the void-galaxy cross-correlation function (VGCF), a statistical measurement of how galaxies are distributed around void centers, which effectively captures the anisotropic features of stacked voids and offers an accurate representation of their average shape. The AP test represents a powerful and well-established method for probing cosmological parameters through the geometric distortions. However, a significant challenge arises from the fact that the AP distortions are not the only ones involved. Indeed, when redshifts are used as proxies for distance, Redshift Space Distortions (RSD), caused by having neglected the contribution of galaxy peculiar velocities to the observed redshift when converting $z$ into distances, also affect the observed structure of voids. These dynamical distortions are degenerate with the AP ones, making it necessary to disentangle the two effects. The impact of these dynamical distortions is typically addressed by modeling together the effect of RSD and the AP on the measured VGCF (see e.g. Nadathur et al. 2020; Hamaus et al. 2020; Woodfinden et al. 2022; Radinović et al. 2023; Hamaus et al. 2022; Verza et al. 2024). However, since the two effects are partly degenerate, we need to model RSD accurately to perform an AP test. To address this problem, several RSD models of increasing complexity have been proposed (see e.g. Paz et al. 2013; Hamaus et al. 2014; Cai et al. 2016; Nadathur & Percival 2018; Hamaus et al. 2022). Although these models were sufficient for cosmological analyses of past datasets, such as BOSS (Dawson et al. 2012), the increased statistics coming with data from the next-generation stage IV spectroscopic redshift surveys such as DESI (DESI Collaboration et al. 2022), Euclid (Euclid Collaboration et al. 2024) and Roman (Spergel et al. 2015; Dore et al. 2019) requires continual improvements in theoretical modeling to match the unprecedented quality of the data. Two approaches are possible. The first is to account for nonlinear effects in the VGCF by either introducing additional nuisance parameters to empirically model them, as in Hamaus et al. (2020), or by incorporating physically motivated nonlinear models directly into the likelihood analysis, as in Nadathur & Percival (2018). The alternative approach is to apply a theoretically justified transformation to the data to minimize the impact of nonlinear effects. In this work, we adopt the latter, commonly known as cosmological reconstruction. More specifically, we use the Zel’dovich approximation (Zel’dovich 1970) to estimate the galaxies’ displacement field from their apparent positions in redshift space using first order Lagrangian perturbation theory. This allows us to remove the peculiar velocity component and restore their real-space positions, effectively eliminating RSDs.
@@
While more sophisticated techniques could be used to account for nonlinear effects (see e.g. Nusser & Branchini 2000; Sarpa et al. 2021; Ganeshaiah Veena et al. 2023; Courtois et al. 2023), we chose here to rely on the Zel’dovich method, as it is generally considered the standard and has proven to be a robust and reliable methodology in galaxy clustering analyses (Padmanabhan et al. 2012; Gil-Mar´ın et al. 2020; DESI Collaboration et al. 2024).
@@
The cosmological reconstruction approach offers two main advantages. First, by removing RSD and their anisotropies, it allows voids to be identified in real space rather than redshift space, significantly reducing the likelihood of selection effects due to being in redshift space (Correa et al. 2022). Second, smaller cosmic voids are particularly susceptible to nonlinear dynamics that cannot be fully captured by linear RSD models. As a result, most RSD-based analyses (see e.g., Hamaus et al. 2020) adopt conservative selection criteria to exclude these smaller voids, which, however, reduces the statistical power of the sample, in particular since small voids are the most abundant. Since the Zel’dovich reconstruction accounts for mildly nonlinear effects, it enables the analysis of smaller-scale voids that are typically excluded in linear RSD analyses, effectively increasing the void sample and reducing statistical errors. In line with this, previous studies such as Nadathur et al. (2019), Woodfinden et al. (2022), and Radinović et al. (2023) have also applied reconstruction techniques, specifically based on the Zel’dovich approximation, to galaxy catalogs prior to void identification, with the aim of mitigating the nonlinearities and selection effects arising in the void-finding process. However, in those studies, the subsequent analysis was performed by correlating void positions in (reconstructed) real space with galaxy positions in redshift space, eventually using RSD models to compare these VGCF measurements with theoretical predictions. This approach overcomes the challenge of identifying voids in redshift space but does not resolve the limitations associated with modeling RSDs in the void profile, with the consequence of not fully disentangling the two effects. Additionally this hybrid approach (voids in reconstructed space and galaxies in redshift space) makes the modelling less straightforward. In this work, we aim at conducting the entire analysis in (reconstructed) real space, cross-correlating void and galaxy positions in reconstructed space. Additionally, we aim to investigate whether this approach can be successfully applied to small-scale voids, thereby increasing both the statistical power of this strategy and the precision of the final constraints, in comparison to the standard analysis performed in redshift space (see e.g. Hamaus et al. 2020, 2022; Verza et al. 2024). Ultimately, in this work we focus on the Zel’dovich reconstruction approach, while a broader comparison with other reconstruction-based methodologies previously mentioned lies beyond the scope of this study and is left for future exploration.
@@ 2 Theoretical background
where $c$ is the speed of light, and $H(z)$ is the Hubble function, which describes the expansion history of the Universe and depends on the assumed cosmological model through the density parameters $\boldsymbol{\Omega}_{i}$ of all the components that contribute to the total energy-mass budget of the Universe, and on the dark energy equation of state. The distance inferred from the redshift may systematically differ from the true distance for two reasons. First, the assumed (or fiducial) cosmological model may differ from the true one. Adopting an incorrect Hubble function $H(z)$, will systematically bias the distance estimate $d(z)$. This results in the introduction of the characteristic Alcock-Pacziński effect, which distorts the 3D mapping of the galaxy distribution. Second, galaxies’ peculiar velocities also contribute to the observed redshift $z_{\mathrm{obs}}$ through the Doppler effect. While smaller in magnitude than the cosmological redshift, neglecting this contribution introduces a systematic bias in the estimated distance, as for the AP test. Moreover, since only the radial component of the peculiar velocity affects the Doppler shift, the inferred mass density field lacks the expected statistical isotropy. Both effects introduce distortions in the 3D mapping from redshift to distance. However, since these distortions are directly linked to cosmological parameters, their detection enables comparison with theoretical models to constrain the matter content $\boldsymbol{\Omega}_{i}$ and the growth rate of structures in the Universe $f(z)$.
@@ 2.1 Alcock-Paczyński distortions
In our analysis, we measure the separation between a galaxy at $(d(z_{1}),\Theta_{1})$ and a cosmic void center at $(d(z_{2}),\Theta_{2})$, both identified in the same spectroscopic sample. Assuming the distant observer approximation and using the void center as the line-of-sight (LOS) (Hamaus et al. 2020), we decompose the void-galaxy separation vector $\mathbf{r}$ into parallel and perpendicular components, such that $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$. Using Eq. (1) and assuming that $H(z_{1})=H(z_{2})$ and $D_{\mathrm{A}}(z_{1})=D_{\mathrm{A}}(z_{2})$ it follows that:
@@
where $\delta z=z_{1}-z_{2}$ and $\delta\theta$ denote the redshift and angular separations of the void-galaxy pair, respectively, and $z$ is the mean redshift of the galaxy sample. This assumes that the size of the sample $l$ is much smaller than its distance to the observer $d(z)\ll l$. $D_{\mathrm{A}}(z)=d(z)/(1+z)$ is the angular diameter distance and the equality holds in a flat Universe. Since $D_{\mathrm{A}}(z)$ depends on $d(z)$, and since $d(z)$ itself is a function of $H(z)$, it follows that $D_{\mathrm{A}}(z)$ also depends on $H(z)$ and thus on the fiducial cosmology.A mismatch between the fiducial and the true cosmology systematically biases pairs separations as follows:
@@
where the primed quantities are evaluated in the fiducial cosmology. In the case of voids, that are standard spheres (Lavaux & Wandelt 2012), meaning that they are spherically symmetric but lacking a characteristic scale as the BAO, the AP test constrains only the product $D_{\mathrm{A}}(z)H(z)$ via the ellipticity parameter $\varepsilon$:
@@
Our aim is to constrain the AP distortion parameter $\varepsilon$ using the multipoles of the VGCF. To enable stacking across voids of varying size, thus focusing on the average profile, we normalize each void-galaxy separation by the void’s effective radius $R$, defined as the cube root of its volume (see Section 3.4). We also account for the AP scaling of $R$ (Hamaus et al. 2020):
@@
Additionally, we express the separation vector $\bf r$} in terms of its absolute value $r$, and its angle with respect to the LOS $\mu=r_{\parallel}/r$. On this new coordinates, the effect of AP distortions is parametrized as (Hamaus et al. 2020):
@@ 2.2 Redshift space distortions
The observed redshift $z_{\mathrm{obs}}$ results from the combination of the cosmological redshift, $z_{\mathrm{c}}$ of a galaxy generated by the Hubble expansion, and a Doppler effect due to the LOS component of its peculiar velocity, $z_{\mathrm{d}}=v_{\parallel}/c$. The observed redshift can be expressed as:
@@
where $f$ is the linear growth rate of the mass density fluctuations, and $\Delta(r)$ the density contrast $\delta$ averaged within a comoving radius $r$:
@@ 2.4 Beyond the linear VGCF model
The range of validity of the linear VGCF model has been widely discussed in recent literature. While Schuster et al. (2023) argue that the linearized continuity equation Eq. (10) remains accurate down to relatively small scales, Nadathur & Percival (2018) highlight its limited applicability in redshift space. These limitations have motivated several improvements to the linear model. For example, Hamaus et al. (2020) introduced nuisance parameters to marginalize over systematic effects, such as nonlinear corrections, spurious voids, and selection biases (see below). Conversely, Nadathur & Percival (2018) incorporated explicit correction terms to account for deviations from linear theory.
@@
Another potential source of disturbance in the model is the assumption of linear galaxy bias, which may not hold in the inner regions of cosmic voids. This issue is thoroughly discussed by Nadathur & Percival (2018); Hamaus et al. (2020) in relation to the VGCF analysis, and additionally by Pollina et al. (2017, 2019); Contarini et al. (2019); Verza et al. (2022). In this work, we adopt a phenomenological approach and use the VGCF model based on an empirical modification of Eqs. (13) and (14), proposed by Hamaus et al. (2022). This model introduces two additional nuisance parameters, $\mathcal{M}$ and $\mathcal{Q}$, compared to the linear RSD model for voids proposed by Cai et al. (2016), and modifies the RSD terms by altering the coefficients that govern the amplitude of the distortions (see Hamaus et al. 2022 for details):
@@
The nuisance parameters $\mathcal{M}$ and $\mathcal{Q}$ are introduced by Hamaus et al. (2020) to account for potential systematics, including potential deviations from the linear theory. This would include deviations in the dynamics and the bias parameter, as well as the potential presence of spurious voids in the sample, which may reduce the amplitude of both the monopole and quadrupole moments of the VGCF (Cousinou et al. 2019). The monopole-like parameter $\mathcal{M}$ regulates the overall amplitude of the VGCF and is intended to account for all factors that may influence the magnitude of the real space void-galaxy clustering. The parameter $\mathcal{Q}$ multiplies the $\mu^{2}$ term, thereby regulating the quadrupole moment. Its role consists in mitigating the loss of amplitude due to spurious voids.
Finally, the model requires the real-space VGCF, $\xi(r)$, along with its volume-averaged counterpart, $\bar{\xi}(r)$, which we estimate phenomenologically from the mean real-space VGCF monopole measured across a suite of 100 mock catalogs (described in Section 3.1). While this choice provides a robust baseline for our analysis, a more data-driven treatment based on alternative prescriptions for modeling $\bar{\xi}(r)$ is left to future work (see e.g. Pisani et al. 2014; Hawken et al. 2017). Furthermore, Eq. (17) already requires the knowledge of the argument $r$ of $\bar{\xi}(r)$. When performing the analysis in redshift space (not in reconstructed space) we need to account for this, therefore we use the method proposed by Hamaus et al. (2020) to evaluate this quantity via iteration using Eq. (14): we start with using $\bar{\xi}(s)$ as initial guess for $\bar{\xi}(r)$, and iteratively calculate $r_{\parallel}$ and $\bar{\xi}(r)$, until convergence is reached.
As with any phenomenological model, it includes uncertainties that primarily affect the amplitude of the VGCF. For this reason, such uncertainties are expected to be absorbed by the parameter $\mathcal{M}$.
We incorporate the AP effect by applying the coordinate transformation described in Eqs. (6) and (7), effectively introducing an explicit dependence of the coordinate $r$, $s$, $\mu$, on the AP distortion parameter $\varepsilon$ in Eq. (17).

### v2

# arXiv:2509.08884v2[astro-ph.CO] 02 Mar 2026

@@
Extracting the Alcock-Paczyński signal from voids:
@@
G. Degni    E. Sarpa,    M. Aubert,    E. Branchini,    A. Pisani,    H.M. Courtois,
(Received 18 September 2025 / Accepted 15 December 2025)
The void-galaxy cross-correlation function is a powerful tool to extract cosmological information. Through the void-galaxy cross-correlation function, cosmic voids – the underdense regions in the galaxy distribution – are used for refined deductions of the Universe’s content by correcting apparent geometric distortions. This study proposes a novel procedure for optimally extracting the Alcock-Paczyǹski (AP) signal from cosmic voids through a cosmological reconstruction technique. Employing cosmological reconstruction, specifically using the Zel’dovich approximation, we estimate the true positions of galaxies from their redshift-space locations, reducing distortions introduced by peculiar velocities. Unlike previous analyses, we identify voids and measure the void-galaxy cross-correlation function directly in reconstructed space. This approach enables us, for the first time, to include in our analysis small nonlinear voids, typically discarded in previous studies, thus enhancing the statistical power of void studies and significantly improving their cosmological constraining power. Reconstruction is particularly effective even at small scales for voids, due to their clean and dynamically simple environment. This ability to recover information encoded on small scales significantly enhances the precision of the analysis, leading to a $\sim 23\%$ improvement in the constraints on the AP parameters compared to previous methods where the analysis is performed in redshift space and, consequently, to a better estimate of the derived cosmological parameters. Our analysis also includes a comprehensive set of consistency checks, demonstrating its robustness. We expect this methodology to yield a substantial gain in constraining power when applied to data from modern large-scale structure surveys.
@@ 1 Introduction
Cosmic voids, vast underdense regions that populate the large-scale structure of the Universe, offer a unique and insightful laboratory for cosmology (Pisani et al. 2019; Kreisch et al. 2019; Moresco et al. 2022). Due to the statistically isotropic nature of structure formation, these initially irregular underdensities evolve toward an approximately spherical shape as matter evacuates their interiors under gravitational dynamics. Although individual voids deviate from sphericity, the process of rescaling by the effective radii and stacking multiple voids yields an average spherically symmetric profile that robustly characterizes their global properties. This averaged spherical structure provides a powerful tool for cosmological analysis. In particular, it enables the application of the Alcock-Paczyǹski (AP) test (Alcock and Paczynski 1979), which relies on the fact that the observed void profile is obtained by converting galaxy positions observed through directly measurable quantities, i.e., angular coordinates ($\mathbf{\Theta}=(\theta,\phi)$), and redshift ($z$), into comoving distances under an assumed cosmological model. If the fiducial model differs from the true one, the inferred galaxy distribution is systematically distorted, breaking the isotropy of stacked voids. By measuring such distortions in the stacked void profile, the AP test allows us to extract valuable information about the expansion history and geometry of the Universe, as well as the dark energy equation of state.
@@
The idea of applying the AP test to cosmic voids was first proposed by Ryden (1995), although its implementation on individual voids proved challenging due to their complex and irregular shapes. This limitation was addressed by Lavaux and Wandelt (2012), who introduced the concept of stacking voids to average out shape irregularities and enhance the statistical power of the test. The first practical applications of the AP test to stacked voids were carried out by Sutter et al. (2012) and Mao et al. (2017). These developments have led to a shift in methodology, favoring the analysis of AP distortions with the void-galaxy cross-correlation function (VGCF) – a statistical measurement of how galaxies are distributed around void centers – which effectively captures the anisotropic features of stacked voids and offers an accurate representation of their average shape. The AP test represents a powerful and well-established method for probing cosmological parameters through the geometric distortions. However, a significant challenge arises from the fact that the AP distortions are not the only ones involved. Indeed, when redshifts are used as proxies for distance, Redshift space distortions (RSD), caused by having neglected the contribution of galaxy peculiar velocities to the observed redshift when converting $z$ into distances, also affect the observed structure of voids. These dynamical distortions are degenerate with the AP ones, making it necessary to disentangle the two effects. The impact of these dynamical distortions is typically addressed by modeling together the effect of RSD and the AP on the measured VGCF (see e.g. Nadathur et al. 2020; Hamaus et al. 2020; Woodfinden et al. 2022; Radinović et al. 2023; Hamaus et al. 2022; Verza et al. 2025). However, since the two effects are partly degenerate, we need to model RSD accurately to perform an AP test. To address this problem, several RSD models of increasing complexity have been proposed (see e.g. Paz et al. 2013; Hamaus et al. 2014; Cai et al. 2016; Nadathur and Percival 2018; Hamaus et al. 2022). Although these models were sufficient for cosmological analyses of past datasets, such as BOSS (Dawson et al. 2012), the increased statistics coming with data from the next-generation stage IV spectroscopic redshift surveys such as DESI (DESI Collaboration et al. 2022), Euclid (Euclid Collaboration et al. 2025), and Roman (Spergel et al. 2015; Dore et al. 2019) requires continual improvements in theoretical modeling to match the unprecedented quality of the data. Two approaches are possible. The first is to account for nonlinear effects in the VGCF by either introducing additional nuisance parameters to empirically model them, as in Hamaus et al. (2020), or by incorporating physically motivated nonlinear models directly into the likelihood analysis, as in Nadathur and Percival (2018). The alternative approach is to apply a theoretically justified transformation to the data to minimize the impact of nonlinear effects. In this work, we adopt the latter, commonly known as cosmological reconstruction. More specifically, we use the Zel’dovich approximation (Zel’dovich 1970) to estimate the galaxies’ displacement field from their apparent positions in redshift space using first order Lagrangian perturbation theory. This allows us to remove the peculiar velocity component and restore their real-space positions, effectively eliminating RSDs.
@@
While more sophisticated techniques could be used to account for nonlinear effects (see e.g. Nusser and Branchini 2000; Sarpa et al. 2021; Ganeshaiah Veena et al. 2023; Courtois et al. 2023), we chose here to rely on the Zel’dovich method, as it is generally considered the standard and has proven to be a robust and reliable methodology in galaxy clustering analyses (Padmanabhan et al. 2012; Gil-Marín et al. 2020; DESI Collaboration. et al. 2025).
@@
The cosmological reconstruction approach offers two main advantages. First, by removing RSD and their anisotropies, it allows voids to be identified in real space rather than redshift space, significantly reducing the likelihood of selection effects due to being in redshift space (Correa et al. 2022). Second, smaller cosmic voids are particularly susceptible to nonlinear dynamics that cannot be fully captured by linear RSD models. As a result, most RSD-based analyses (see e.g. Hamaus et al. 2020) adopt conservative selection criteria to exclude these smaller voids, which, however, reduces the statistical power of the sample, in particular since small voids are the most abundant. Since the Zel’dovich reconstruction accounts for mildly nonlinear effects, it enables the analysis of smaller-scale voids that are typically excluded in linear RSD analyses, effectively increasing the void sample and reducing statistical errors. In line with this, previous studies such as Nadathur et al. (2019), Woodfinden et al. (2022), and Radinović et al. (2023) have also applied reconstruction techniques, specifically based on the Zel’dovich approximation, to galaxy catalogs prior to void identification, with the aim of mitigating the nonlinearities and selection effects arising in the void-finding process. However, in those studies, the subsequent analysis was performed by correlating void positions in (reconstructed) real space with galaxy positions in redshift space, eventually using RSD models to compare these VGCF measurements with theoretical predictions. This approach overcomes the challenge of identifying voids in redshift space but does not resolve the limitations associated with modeling RSDs in the void profile, with the consequence of not fully disentangling the two effects. Additionally, this hybrid approach (voids in reconstructed space and galaxies in redshift space) makes the modeling less straightforward. In this work, we aim at conducting the entire analysis in (reconstructed) real space, cross-correlating void and galaxy positions in reconstructed space. Additionally, we aim to investigate whether this approach can be successfully applied to small-scale voids, thereby increasing both the statistical power of this strategy and the precision of the final constraints, in comparison to the standard analysis performed in redshift space (see e.g. Hamaus et al. 2020, 2022; Verza et al. 2025). Ultimately, in this work we focus on the Zel’dovich reconstruction approach, while a broader comparison with other reconstruction-based methodologies previously mentioned lies beyond the scope of this study and is left for future exploration.
@@ 2 Theoretical background
where $c$ is the speed of light, and $H(z)$ is the Hubble function, which describes the expansion history of the Universe and depends on the assumed cosmological model through the density parameters $\boldsymbol{\Omega}_{i}$ of all the components that contribute to the total energy-mass budget of the Universe, and on the dark energy equation of state. The distance inferred from the redshift may systematically differ from the true distance for two reasons. First, the assumed (or fiducial) cosmological model may differ from the true one. Adopting an incorrect Hubble function $H(z)$ will systematically bias the distance estimate $d(z)$. This results in the introduction of the characteristic Alcock-Pacziński effect, which distorts the 3D mapping of the galaxy distribution. Second, galaxies’ peculiar velocities also contribute to the observed redshift $z_{\mathrm{obs}}$ through the Doppler effect. While smaller in magnitude than the cosmological redshift, neglecting this contribution introduces a systematic bias in the estimated distance, as for the AP test. Moreover, only the radial component of the peculiar velocity affects the Doppler shift, the inferred mass density field therefore lacks the expected statistical isotropy. Both effects introduce distortions in the 3D mapping from redshift to distance. However, since these distortions are directly linked to cosmological parameters, their detection enables comparison with theoretical models to constrain the matter content $\boldsymbol{\Omega}_{i}$ and the growth rate of structures in the Universe $f(z)$.
@@ 2.1 Alcock-Paczyński distortions
In our analysis, we measure the separation between a galaxy at $(d(z_{1}),\Theta_{1})$ and a cosmic void center at $(d(z_{2}),\Theta_{2})$, both identified in the same spectroscopic sample. Assuming the distant observer approximation and using the void center as the line-of-sight (LOS) (Hamaus et al. 2020), we decompose the void-galaxy separation vector $\mathbf{r}$ into parallel and perpendicular components, such that $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$. Using Eq. (1), and assuming that $H(z_{1})=H(z_{2})$ and $D_{\mathrm{A}}(z_{1})=D_{\mathrm{A}}(z_{2})$, it follows that
@@
where $\delta z=z_{1}-z_{2}$ and $\delta\theta$ denote the redshift and angular separations of the void-galaxy pair, respectively, and $z$ is the mean redshift of the galaxy sample. This assumes that the size of the sample $l$ is much smaller than its distance to the observer $d(z)\ll l$. $D_{\mathrm{A}}(z)=d(z)/(1+z)$ is the angular diameter distance and the equality holds in a flat Universe. Since $D_{\mathrm{A}}(z)$ depends on $d(z)$, and since $d(z)$ itself is a function of $H(z)$, it follows that $D_{\mathrm{A}}(z)$ also depends on $H(z)$ and thus on the fiducial cosmology. A mismatch between the fiducial and the true cosmology systematically biases pairs separations as follows
@@
where the primed quantities are evaluated in the fiducial cosmology. In the case of voids, that are standard spheres (Lavaux and Wandelt 2012), meaning that they are spherically symmetric but lacking a characteristic scale as the BAO, the AP test constrains only the product $D_{\mathrm{A}}(z)H(z)$ via the ellipticity parameter $\varepsilon$:
@@
Our aim is to constrain the AP distortion parameter $\varepsilon$ using the multipoles of the VGCF. To enable stacking across voids of varying size, thus focusing on the average profile, we normalize each void-galaxy separation by the void’s effective radius $R$, defined as the cube root of its volume (see Section 3.4). We also account for the AP scaling of $R$ (Hamaus et al. 2020)
@@
Additionally, we express the separation vector $\bf r$} in terms of its absolute value $r$, and the cosine of its angle with respect to the LOS $\mu=r_{\parallel}/r$. In these new coordinates, the effect of AP distortions is parametrized as (Hamaus et al. 2020)
@@ 2.2 Redshift space distortions
The observed redshift $z_{\mathrm{obs}}$ results from the combination of the cosmological redshift, $z_{\mathrm{c}}$ of a galaxy generated by the Hubble expansion, and a Doppler effect due to the LOS component of its peculiar velocity, $z_{\mathrm{d}}=v_{\parallel}/c$. The observed redshift can be expressed as
@@
where $f$ is the linear growth rate of the mass density fluctuations, and $\Delta(r)$ the density contrast $\delta$ averaged within a comoving radius $r$
@@ 2.4 Beyond the linear VGCF model
The range of validity of the linear VGCF model has been widely discussed in recent literature. While Schuster et al. (2023) argue that the linearized continuity equation Eq. (10) remains accurate down to relatively small scales, Nadathur and Percival (2018) highlight its limited applicability in redshift space. These limitations have motivated several improvements to the linear model. For example, Hamaus et al. (2020) introduced nuisance parameters to marginalize over systematic effects, such as nonlinear corrections, spurious voids, and selection biases (see below). Conversely, Nadathur and Percival (2018) incorporated explicit correction terms to account for deviations from linear theory.
@@
Another potential source of disturbance in the model is the assumption of linear galaxy bias, which may not hold in the inner regions of cosmic voids. This issue is thoroughly discussed by Nadathur and Percival (2018); Hamaus et al. (2020) in relation to the VGCF analysis, and additionally by Pollina et al. (2017, 2019); Contarini et al. (2019); Verza et al. (2022). In this work, we adopt a phenomenological approach and use the VGCF model based on an empirical modification of Eqs. (13) and (14), proposed by Hamaus et al. (2022). This model introduces two additional nuisance parameters, $\mathcal{M}$ and $\mathcal{Q}$, compared to the linear RSD model for voids proposed by Cai et al. (2016), and modifies the RSD terms by altering the coefficients that govern the amplitude of the distortions (see Hamaus et al. 2022 for details):
@@
The nuisance parameters $\mathcal{M}$ and $\mathcal{Q}$ are introduced by Hamaus et al. (2020) to account for potential systematics, including potential deviations from the linear theory. This would include deviations in the dynamics and the bias parameter, as well as the potential presence of spurious voids in the sample, which may reduce the amplitude of both the monopole and quadrupole moments of the VGCF (Cousinou et al. 2019). The monopole-like parameter $\mathcal{M}$ regulates the overall amplitude of the VGCF and is intended to account for all factors that may influence the magnitude of the real-space void-galaxy clustering. The parameter $\mathcal{Q}$ multiplies the $\mu^{2}$ term, thereby regulating the quadrupole moment. Its role consists of mitigating the loss of amplitude due to spurious voids.
Finally, the model requires the real-space VGCF, $\xi(r)$, along with its volume-averaged counterpart, $\bar{\xi}(r)$, which we estimate phenomenologically from the mean real-space VGCF monopole measured across a suite of 100 mock catalogs (described in Section 3.1). While this choice provides a robust baseline for our analysis, a more data-driven treatment based on alternative prescriptions for modeling $\bar{\xi}(r)$ is left to future work (see e.g. Pisani et al. 2014; Hawken et al. 2017). Furthermore, Eq. (17) already requires the knowledge of the argument $r$ of $\bar{\xi}(r)$. When performing the analysis in redshift space (not in reconstructed space) we need to account for this, therefore we use the method proposed by Hamaus et al. (2020) to evaluate this quantity via iteration using Eq. (14): we start with using $\bar{\xi}(s)$ as initial guess for $\bar{\xi}(r)$, and iteratively calculate $r_{\parallel}$ and $\bar{\xi}(r)$, until convergence is reached. As with any phenomenological model, it includes uncertainties that primarily affect the amplitude of the VGCF. For this reason, such uncertainties are expected to be absorbed by the parameter $\mathcal{M}$. We incorporate the AP effect by applying the coordinate transformation described in Eqs. (6) and (7), effectively introducing an explicit dependence of the coordinate $r$, $s$, $\mu$, on the AP distortion parameter $\varepsilon$ in Eq. (17).